OpenStudy (anonymous):

MEDAL WILL BE AWARDED AND I WILL BECOME YOUR FAN Choose the correct simplification of the expression (3x)4. 81x4 12x5 81x5 12x4

OpenStudy (anonymous):

@JazzyFrizo16 @Helpmeplease99 @mathstudent55

OpenStudy (anonymous):

is it (3x)^4

OpenStudy (anonymous):

yup

OpenStudy (johnweldon1993):

\[(3x)^4 = 3^4 \times x^4\] so what is 3^4?

OpenStudy (anonymous):

3*3*3*3=81

OpenStudy (johnweldon1993):

Right... so \[81 \times x^4 = 81x^4\] and that would be your simplification

OpenStudy (anonymous):

i agree with john.

OpenStudy (anonymous):

Wow thank you i was about to get it wrong

OpenStudy (anonymous):

And thank you please99

OpenStudy (johnweldon1993):

No problem!

OpenStudy (anonymous):

:)